Harry started a new job in 2003. His salary was $31,000. At the beginning of the next year he will receive a raise of $1500. Assume he will receive the same raise every year.
a. Write a function rule for finding Harry's salary after 2003.
b. Find Harry's salary in 2008.

Answer by misscrt(37)

Harry started a new job in 2003. His salary was $31,000. At the beginning of the next year he will receive a raise of $1500. Assume he will receive the same raise every year.
________________________________________________________________________________
a)
S=31000+1500Y
Where S is salary and Y is # of years
b)
Salary in 5 years;
=31000+1500(5)
=31000+7500
=38500
Hope that helps.
Greetings,
misscrt